Please help me solve the following system of equation.
-5x-3y=12
-2x-5y=1
x=
y=
Respectfully,
M. Rowser

Answer by ewatrrr(24785)

-5x-3y=12
-2x-5y=1 Multiply by -5/2 and ADD EQs
(19/2)y = 19/2, y = 1 and x = -3 +-2x+=+6